Use the method disks/rings to determine the volume of the solid obtained by rotating the region bounded by the given curves about the given axis of y=2x+1, x=4 and y= 3 about the line x=-4.


1
Expert's answer
2021-04-14T02:43:55-0400

The given lines are represented as-




The region formed by rotating the above regions is-



So, the inner and outer radii are as follows:


Inner Radius "r=4+\dfrac{1}{2}(y-1)=\dfrac{y}{2}+\dfrac{7}{2}"


Outer radius "R=4+4=8"


Area of the ring is

"A(x)=\pi [R^2-r^2]"


   "=\pi[(8)^2-(\dfrac{y}{2}+\dfrac{7}{2})^2]"


   "=\pi(\dfrac{207}{4}-\dfrac{7}{2}y-\dfrac{y^2}{4})"



Then the volume of the solid obtained is


 "V=\int_3^9A(x)dy=\pi(\dfrac{207}{4}-\dfrac{7}{2}y-\dfrac{y^2}{4})=\int_3^9 \pi(\dfrac{207}{4}-\dfrac{7}{2}y-\dfrac{y^2}{4})dy"


  "=\pi(\dfrac{207y}{4}-\dfrac{7y^2}{4}-\dfrac{y^3}{12})|_3^9=126\pi"